OpenStudy (anonymous):

What are the solution(s) to the following system of equations? y = –x2 y = –x

OpenStudy (anonymous):

y = –x2 y = –x

OpenStudy (cwrw238):

-x^2 = -x x - x^2 = 0 x(1-x) = 0 x = 0 or 1

OpenStudy (anonymous):

so which one would it be a.(1, –1) and (–1, –1) b. (0, 0) c.(0, 0) and (1, –1) d.(0, 0) and (–1, 1)

OpenStudy (cwrw238):

oh right - sorry - i didnt calculate the value of y when x = 0, y = 0 when x = 1 , y = -1 so the answer is (0,0) and (1,-1) c

OpenStudy (cwrw238):

i got these from plugging in values of x into y = -x

OpenStudy (cwrw238):

happy with that?

OpenStudy (anonymous):

thank you :)

OpenStudy (cwrw238):

yw